# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

### v1

@@ 7.3.2. Synthetic Data
An important question in the field of Language Models is that of synthetic data (see e.g. [WWZ24, LWLZD24]) for LLM training. An interesting question is to determine how much value can be learned by pre-training LLMs on synthetic data coming from gameplay on Xent Games. The long-context window and other challenging gameplay aspects could in principle lead pre-trained models to learn deeper and more nuanced representations of information, enabling them to capture subtle notions associated with the implicit knowledge tasks.
@@ 7.3.4. Putting Things Together
In principle, these models can all be the same, and they can evolve over time. If one leverages Xent Games as means to train new models (as suggested in Sections 7.3.2 and 7.3.3 above), then this can lead to a self-improvement loop of LLMs, which is an interesting alley to pursue: this is one way in which open-endedness ideas could (in principle) lead to an augmentation of capabilities for LLMs in new dimensions.

### v2

@@ 7.3.3. Synthetic Data
An important question in the field of Language Models is that of synthetic data (see e.g. [WWZ24, LWLZD24]) for LLM training. An interesting question is to determine how much value can be learned by pre-training LLMs on synthetic data coming from gameplay of Xent Games. The long-context window and other challenging gameplay aspects could in principle lead pre-trained models to learn deeper and more nuanced representations of information, enabling them to capture subtle notions associated with implicit knowledge tasks.
@@ 7.3.5. Putting Things Together
In principle, these models can all be the same, and they can evolve over time. If one leverages Xent Games as means to train new models (as suggested in Sections 7.3.3 and 7.3.4 above), then this can lead to a self-improvement loop of LLMs, which is an interesting alley to pursue: this is one way in which evolutionary ideas could (in principle) lead to an augmentation of capabilities for LLMs in new, open-ended dimensions.